# Deep Networks Always Grok and Here is Why

### v1

@@ Abstract
Grokking, or delayed generalization, is a phenomenon where generalization in a deep neural network (DNN) occurs long after achieving near zero training error. Previous studies have reported the occurrence of grokking in specific controlled settings, such as DNNs initialized with large-norm parameters or transformers trained on algorithmic datasets. We demonstrate that grokking is actually much more widespread and materializes in a wide range of practical settings, such as training of a convolutional neural network (CNN) on CIFAR10 or a Resnet on Imagenette. We introduce the new concept of delayed robustness, whereby a DNN groks adversarial examples and becomes robust, long after interpolation and/or generalization. We develop an analytical explanation for the emergence of both delayed generalization and delayed robustness based on a new measure of the local complexity of a DNN’s input-output mapping. Our local complexity measures the density of the so-called “linear regions” (aka, spline partition regions) that tile the DNN input space, and serves as a utile progress measure for training. We provide the first evidence that for classification problems, the linear regions undergo a phase transition during training whereafter they migrate away from the training samples (making the DNN mapping smoother there) and towards the decision boundary (making the DNN mapping less smooth there). Grokking occurs post phase transition as a robust partition of the input space emerges thanks to the linearization of the DNN mapping around the training points. bit.ly/grok-adversarial.
@@
Figure 1: Deep Neural Networks grok robustness. When training a ResNet18 on CIFAR10, without any controlled initialization as in Liu et al. (2022), the network starts grokking adversarial examples generated using Projected Gradient Descent (Madry et al., 2017) after $10^{4}$ optimization steps (top) and attains almost equal robustness and generalization performance after $2\times 10^{5}$ steps. We see that, prior to grokking, the network undergoes a phase change during training in the local complexity, i.e., the local density of non-linearities in the input space (bottom). After test accuracy converges, the network starts migrating its non-linearities away from the data points and closer to the decision boundary (see Figure 2), eventually reducing the complexity of the learned function around the data points. This increase and subsequent decrease in local non-linearity is a phenomenon visible for a wide variety of networks and training settings (see Footnote 2). In this paper, we show that this particular training dynamic always results in delayed generalizaion or robustness.
@@ 1 Introduction
Figure 2: Emergence of Robust Partition. We train a 4-layer ReLU MLP of $200$ width, on $1K$ samples from MNIST for $10^{5}$ optimization steps, with batch size $200$. We see that the network starts grokking adversarial examples after approximately $10^{4}$ optimization steps (top-left). The local complexity around data points (bottom-left) follows a double descent curve with the final descent starting approximately after $10^{4}$ optimization steps as well. Where do the non-lienarities migrate to? In the middle and right images we present analytically computed visualizations of the DNN input space partition (Humayun et al., 2023a). The partition or linear regions are visualized across a 2D domain in the input space, that intersects three training samples. We see that during the final descent in local complexity, a unique structure emerges in the DNN partition geometry, where a large number of non-linearities (black lines) have concentrated around the decision boundary (red line). We dub this phenomenon Region Migration. Animation for the whole training run in bit.ly/grok-splinecam .
@@
Question. How subjective is the onset of grokking on the test data? When grokking does not manifest as a measurable change in the test set performance, could there exist an alternate test dataset for which grokking would occur?
@@
Observation. For a number of training settings, with standard initialization with or without weight decay, DNNs grok adversarial samples long after generalizing on the test dataset. We dub this phenomenon delayed robustness, a novel form of grokking previously unreported.
We make this observation for a number of training settings including for fully connected networks trained on MNIST (Figure 2), CNNs trained on CIFAR10 and CIFAR100 (Footnote 2), and ResNet18 without batch-normalization, trained on CIFAR10 (Figure 1) and Imagenette (Footnote 2). We generate adversarial examples after each training step using $\ell_{\infty}$-PGD with varying $\epsilon\in\{0.06,0.10,0.13,0.16,0.20\}$, $\alpha=0.0156$ and $10$ ($100$ for MNIST) PGD steps. This observation answers our initial question: indeed there can exist a dataset other than the test dataset for which grokking manifests in classification accuracy. Moreover, we observe that the same phenomenon occurs when test set grokking is induced via initialization scaling (Figure 7).
@@
Figure 3: Curvature and complexity. Visual depiction of Equation 2 with a toy affine spline $S:\mathbb{R}^{2}\rightarrow\mathbb{R}$, obtained by training an MLP to regress the piecewise function $f(x_{1},x_{2})=\{\sin(x_{1})+\cos(x_{2})\}\mathbbm{1}_{x_{1}<0}$. Regions in the input space partition $\Omega$ (left) and the graph of the affine spline function (right) are randomly colored. The spline partition has significantly higher density of non-linearities for $x_{1}<0$, i.e., the local complexity is higher where the function has more curvature.
It has previously been established that both robustness and generalization are a function of the expressivity (Xu & Mannor, 2012; Li et al., 2022) as well as the local linearity (Qin et al., 2019; Balestriero & LeCun, 2023; Humayun et al., 2023b) of a DNN. To explain grokking, we propose a novel complexity measure based on the local non-linearity of the DNN. Our novel measure does not rely on the dataset, labels, or loss function that is used during training. It behaves as a progress measure (Barak et al., 2022; Nanda et al., 2023) that exhibits dynamics correlating with the onset of both delayed generalization and robustness, opening new avenues to study grokking and DNN training dynamics. We show that DNNs undergo phase a change in the local complexity (LC) averaged over data points. Based on these dynamics, we come to the following conclusion:
Claim. Grokking occurs due to the emergence of a robust input space partition by a DNN, through a linearization of the DNN function around training points as a consequence of training dynamics.
We summarize the contributions as follows:
  • We observe for the first time delayed robustness, a novel form of grokking for DNNs that occurs for a whide range of training settings and co-occurs with delayed generalization.
  • We develop a novel progress measure (Barak et al., 2022) for DNN’s based on the local complexity of a DNN’s input space partition. Our proposed measure is a proxy for the DNN’s expressivity, it is task agnostic yet informative of training dynamics. Using our measure, we detect three phases in training: two descent phases and an ascent phase. This is the first time that such dynamics in a DNN’s partition are reported. We crucially observe that a DNN’s partition regions concentrate around the decision boundary long after interpolation, a phenomenon we term as region migration.
@@
We organize the rest of the paper as follows. In Section. 2 we overview the spline interpretation of deep networks and introduce our proposed local complexity measure. We also draw contrasts with common interpretability frameworks, e.g., the commonly used notion of circuits (Olah et al., 2020) in mechanistic interpretability. In Section. 3 we introduce the double descent characteristics of local complexity and connect region migration, i.e., the final phase of the double descent LC dynamics with grokking. We also present results showing that grokking does not happen when using batch normalization and provide theoretical justification. We present results connecting grokking with parameterization and memorization. Finally we mention conclusions drawn from the presented results and limitations of our analysis.
@@ 2.1 Deep Networks are Affine Spline Operators
Formulations like that in Equation 2 that represent DNNs as continuous piecewise affine splines, have previously been employed to make theoretical studies amenable to actual DNNs, e.g. in generative modeling (Humayun et al., 2022), network pruning (You et al., 2021), and OOD detection (Ji et al., 2022). Empirical estimates of the density of linear regions in the spline partition have also been employed in sensitivity analysis (Novak et al., 2018), quantifying non-linearity (Gamba et al., 2022), quantifying expressivity (Raghu et al., 2017) or to estimate the complexity of spline functions (Hanin & Rolnick, 2019). We demonstrate the relationship between function curvature and linear region density through a toy example in Figure 3. In Figure 3-left and Figure 2-(middle,right), any contiguous line is a non-linearity in the input space, corresponding to a single neuron of the network. All the non-linearities re-orient themselves during training to be able to obtain the target function (Figure 3-right). Therefore, in Figure 3, we see that DNN partitions have higher density of linear regions/non-linearities/knots in the spline partition, where the function curvature is higher.
@@ 2.2 Measuring Local Complexity using the Deep Network Spline Partition
Figure 5: Deformation with depth. Change of average eccentricity (Xu et al., 2021) of the input space neighborhoods ${\bm{V}}_{x}$ by different layers of a CNN trained on the CIFAR10 dataset, for different radius $r$. We see that, for larger radius, the deformation increases with depth almost exponentially. For $r\leq 0.014$ deformation is low, indicating that smaller radius neighborhoods are reliable for LC computation on deeper networks. Values are averaged over neighborhoods sampled for $1000$ training points from CIFAR10. For ResNet18 see Figure 16.
@@
is the number of hyperplanes from layer $\ell$ intersecting $\Phi$. We consider $\mathcal{N}$ as a proxy for local complexity for any neighborhood $\Phi$. To make computation tractable, let, $\Phi\approx\widehat{\Phi}=conv(f_{1:\ell-1}({\bm{V}}))$. Therefore, for $\widehat{\Phi}$, any sign changes in layer $\ell$ pre-activations is due to the corresponding neuron hyperplanes intersecting $conv({\bm{V}})$. Therefore for a single layer, the local complexity (LC) for a sample in the input space can be approximated by the number of neuron hyperplanes that intersect $\bm{V}$ embedded to that layers input space. If we consider input space neighborhoods with the same volume, then local complexity measures the un-normalized density of non-linearity in an input space locality. We highlight that this is tied to the VC-dimension of (ReLU) DNN (Bartlett et al., 2019) where the more regions are present the more expressive the decision boundary can be (Montufar et al., 2014). In Figure 4, we provide a visual explanation of our method for local complexity approximation through a cartoon schematic diagram. To summarize, we consider randomly oriented $P$ dimensional $\ell_{1}$ norm balls with radius $r$, i.e., cross-polytopes centered on any given data point $x$ as a frame defining the neighborhood. We therefore follow the steps entailed in Figure 4 in a layerwise fashion, to approximate the local complexity in the prescribed neighborhood.
@@
Refer to caption Optimization Steps
Refer to captionRefer to captionRefer to caption Opt. Step 77035                 Opt. Step 83375                 Opt. Step 95381
Figure 7: Grokking visualized. We induce grokking by randomly initializing a $4$ depth $200$ width ReLU MLP and scaling the initialized parameters by eigth following (Liu et al., 2022). In the leftmost figure, we can see that the grokking is visible for both the test samples as well as adversarial examples generated using the test set. We see that the network robustness, periodically increases. By visualization the partition and curvature of the function across a 2D slice of the input space (Humayun et al., 2023a), we see that the network periodically increases the concentration of non-linearity around its decision boundary, making the boundary sharper at each robustness peak. This occurs even when the network doesn’t undergo delayed generalization (Figure 2). As the local complexity around the decision boundary increases, the local complexity around data points farther from the decision boundary decreases (Figure 18).
@@
Figure 8: Region migration in nodular addition. By measuring the local complexity for the GeLU activated fully connected layers of a Transformer architecture, we see that here as well, region migration occurs during grokking.
Sensitivity of approximation to $P$ and $r$
One of the possible limitations of local complexity measure is the deformation of the local neighborhood when its passed through a network from layer to layer, as shown in Figure 4. For different radius $r$ of the input space neighborhood ${\bm{V}}_{x}$ centered on any arbitrary data point $x$, we compute the change of graph eccentricity (Xu et al., 2021) by different layers of a CNN to measure the degree of deformation by each layer. We present the results in Figure 5 for $1000$ different training data points for a CNN trained on CIFAR10. The higher the deformation, the less reliable the approximation. Here, layer index 00 corresponds to the input space. We see that below a certain radius value, deformation by the CNN is limited and does not exponentially increase. In subsequent experiments, e.g., Figure 19, we have also observed that the dynamics of local complexity is similar between large and small $r$ neighborhoods. We present more validation experiments in Appendix A.
Experimental Setup
@@ 3 Local Complexity Training Dynamics and Grokking
3.1 Emergence of a Robust Partition.
@@
Local complexity as progress measure
@@
Relationship with Circuits

### v2

@@ Abstract
Grokking, or delayed generalization, is a phenomenon where generalization in a deep neural network (DNN) occurs long after achieving near zero training error. Previous studies have reported the occurrence of grokking in specific controlled settings, such as DNNs initialized with large-norm parameters or transformers trained on algorithmic datasets. We demonstrate that grokking is actually much more widespread and materializes in a wide range of practical settings, such as training of a convolutional neural network (CNN) on CIFAR10 or a Resnet on Imagenette. We introduce the new concept of delayed robustness, whereby a DNN groks adversarial examples and becomes robust, long after interpolation and/or generalization. We develop an analytical explanation for the emergence of both delayed generalization and delayed robustness based on the local complexity of a DNN’s input-output mapping. Our local complexity measures the density of so-called “linear regions” (aka, spline partition regions) that tile the DNN input space and serves as a utile progress measure for training. We provide the first evidence that, for classification problems, the linear regions undergo a phase transition during training whereafter they migrate away from the training samples (making the DNN mapping smoother there) and towards the decision boundary (making the DNN mapping less smooth there). Grokking occurs post phase transition as a robust partition of the input space thanks to the linearization of the DNN mapping around the training points. Web: bit.ly/grok-adversarial.
@@
Figure 1: Deep Neural Networks grok robustness. When training a ResNet18 on CIFAR10, without any controlled initialization as in Liu et al. (2022), the network starts grokking adversarial examples generated using Projected Gradient Descent (Madry et al., 2017) after $10^{4}$ optimization steps (top) and attains almost equal robustness and generalization performance after $2\times 10^{5}$ steps. We see that, prior to grokking, the network undergoes a phase change during training in the local complexity, i.e., the local density of spline partition regions in the input space (bottom). After test accuracy converges, the network starts migrating its non-linearities away from the data points and closer to the decision boundary (see Figure 2), eventually reducing the complexity of the learned function around the data points. This increase and subsequent decrease in local non-linearity is a phenomenon visible for a wide variety of networks and training settings (see Footnote 2). In this paper, we show that this particular training dynamic always results in delayed generalization or robustness.
@@ 1 Introduction
Figure 2: Emergence of Robust Partition. We train a 4-layer ReLU Multi Layer Perceptron (MLP) of $200$ width, on $1K$ samples from MNIST for $10^{5}$ optimization steps, with batch size $200$. We see that the network starts grokking adversarial examples after approximately $10^{4}$ optimization steps (top-left). The local complexity around data points (bottom-left) follows a double descent curve with the final descent starting approximately after $10^{4}$ optimization steps as well. Where do the non-linearities migrate to? In the middle and right images we present analytically computed visualizations of the DNN input space partition (Humayun et al., 2023a). The partition or linear regions are visualized across a 2D domain in the input space, that intersects three training samples. We see that during the final descent in local complexity, a unique structure emerges in the DNN partition geometry, where a large number of non-linearities (black lines) therefore linear regions, have concentrated around the decision boundary (red line). We dub this phenomenon Region Migration. Animation for an entire training run in bit.ly/grok-splinecam.
@@
Question. How subjective is the onset of grokking on the test data? When grokking does not manifest as a measurable change in the test set performance, could there exist an alternate set of samples for which grokking would occur?
@@
Observation. For a number of training settings, with standard initialization with or without weight decay, DNNs grok adversarial samples long after generalizing on the test dataset. We dub this novel, previously unreported form of grokking delayed robustness.
We make this observation for a number of training settings including for fully connected networks trained on MNIST (Figure 2), Convolutional Neural Networks (CNNs) trained on CIFAR10 and CIFAR100 (Footnote 2), ResNet18 without batch-normalization, trained on CIFAR10 (Figure 1) and Imagenette (Footnote 2), and a GPT-based Architecture trained on Shakespeare Text (Figure 9). We generate adversarial examples after each training step using $\ell_{\infty}$-PGD with varying $\epsilon\in\{0.03,0.06,0.10,0.13,0.16,0.20\}$, $\alpha=0.0156$ and $10$ ($100$ for MNIST) PGD steps. This observation answers our initial question: indeed there can exist a dataset other than the test dataset for which grokking manifests in classification accuracy. Moreover, we observe that the same phenomenon occurs when test set grokking is induced via initialization scaling (Figure 7) or when training transformers on Modular Addition (Figure 8).
@@
Figure 3: Curvature and complexity. Visual depiction of Equation 2 with a toy affine spline $S:\mathbb{R}^{2}\rightarrow\mathbb{R}$, obtained by training an MLP to regress the piecewise function $f(x_{1},x_{2})=\{\sin(x_{1})+\cos(x_{2})\}\mathbbm{1}_{x_{1}<0}$. Regions in the input space partition $\Omega$ (left) and the graph of the affine spline function (right) are randomly colored. The spline partition has significantly higher density of non-linearities for $x_{1}<0$, i.e., the local complexity is higher where the learned function has more curvature.
It has previously been established that both robustness and generalization are a function of the expressivity (Xu & Mannor, 2012; Li et al., 2022) as well as the local linearity (Qin et al., 2019; Balestriero & LeCun, 2023; Humayun et al., 2023c) of a DNN. To explain grokking, we propose a novel complexity measure based on the local non-linearity of the DNN. Our novel measure does not rely on the dataset, labels, or loss function that is used during training. It behaves as a progress measure (Barak et al., 2022; Nanda et al., 2023) that exhibits dynamics correlating with the onset of both delayed generalization and robustness, opening new avenues to study grokking and DNN training dynamics. We show that DNNs undergo a phase change in the local complexity (LC) averaged over data points. Based on these dynamics, we come to the following conclusion:
Claim. Grokking occurs due to the emergence of a robust input space partition by a DNN, through a linearization of the DNN function around training points as a consequence of the training dynamics. This leads to larger linear regions around training points, and accumulation of non-linearities/linear regions around the decision boundary.
We summarize our contributions as follows:
  • We observe for the first time delayed robustness, a novel form of grokking for DNNs that occurs for a wide range of training settings and co-occurs with delayed generalization.
  • We develop a novel progress measure (Barak et al., 2022) for DNN’s based on the local complexity of a DNN’s input space partition. Our proposed measure is a proxy for the DNN’s expressivity, it is task agnostic yet informative of training dynamics. Using our measure, we detect three phases in training: two descent phases and an ascent phase. This is the first time that such dynamics in a DNN’s partition are reported. We crucially observe that a DNN’s partition regions concentrate around the decision boundary long after interpolation, a phenomenon we term region migration.
@@
We organize the rest of the paper as follows. In Section. 2 we overview the spline interpretation of deep networks and introduce our proposed local complexity measure. We also draw contrasts with common interpretability frameworks, e.g., the commonly used notion of circuits (Olah et al., 2020) in mechanistic interpretability. In Section. 3 we introduce the double descent characteristics of local complexity and connect region migration, i.e., the final phase of the double descent LC dynamics with grokking. We also present results showing that grokking does not happen when using batch normalization and provide theoretical justification. We present results connecting grokking with parameterization and memorization. Finally we draw conclusions from our results and discuss the limitations of our analysis.
@@ 2.1 Deep Networks are Affine Spline Operators
Formulations like that in Equation 2 that represent DNNs as continuous piecewise affine splines, have previously been employed to make theoretical studies amenable to actual DNNs, e.g. in generative modeling (Humayun et al., 2022), network pruning (You et al., 2021), and OOD detection (Ji et al., 2022). Empirical estimates of the density of linear regions in the spline partition have also been employed in sensitivity analysis (Novak et al., 2018), quantifying non-linearity (Gamba et al., 2022), quantifying expressivity (Raghu et al., 2017) or to estimate the complexity of spline functions (Hanin & Rolnick, 2019). We demonstrate the relationship between function curvature and linear region density through a toy example in Figure 3. In Figure 3-left and Figure 2-(middle,right), any contiguous line is a non-linearity in the input space, corresponding to a single neuron of the network. All the non-linearities re-orient themselves during training to be able to obtain the target function (Figure 3-right). Therefore, in Figure 3, we see that DNN partitions have higher density of linear regions/non-linearities/knots in the spline partition, where the target function curvature is non-zero.
@@ 2.2 Measuring Local Complexity using the Deep Network Spline Partition
Figure 5: Deformation with depth. Change of average eccentricity (Xu et al., 2021) of the input space neighborhoods ${\bm{V}}_{x}$ by different layers of a CNN trained on the CIFAR10 dataset, for different radius $r$. We see that, for larger radius, the deformation increases with depth almost exponentially. For $r\leq 0.014$ deformation is low, indicating that smaller radius neighborhoods are reliable for LC computation on deeper networks. Values are averaged over neighborhoods sampled for $1000$ training points from CIFAR10. For ResNet18, see Figure 23.
@@
is the number of hyperplanes from layer $\ell$ intersecting $\Phi$. We consider $\mathcal{N}$ as a proxy for the local complexity of any neighborhood $\Phi$. To make computation tractable, let, $\Phi\approx\widehat{\Phi}=conv(f_{1:\ell-1}({\bm{V}}))$. Therefore, for $\widehat{\Phi}$, any sign changes in layer $\ell$ pre-activations is due to the corresponding neuron hyperplanes intersecting $conv({\bm{V}})$. For a single layer, the local complexity (LC) for a sample in the input space can be approximated by the number of neuron hyperplanes that intersect $\bm{V}$ embedded to that layers input space. If we consider input space neighborhoods with the same volume, then our approximation method measures the un-normalized density of non-linearity in an input space locality, which we consider a proxy for local complexity. We highlight that this is tied to the VC-dimension of (ReLU) DNN (Bartlett et al., 2019) where the more regions are present the more expressive the decision boundary can be (Montufar et al., 2014). In Figure 4, we provide a visual explanation of our method for local complexity approximation through a cartoon schematic diagram. To summarize, we consider randomly oriented $P$ dimensional $\ell_{1}$ norm balls with radius $r$, i.e., cross-polytopes centered on any given data point $x$ as a frame defining the neighborhood. We therefore follow the steps entailed in Figure 4 in a layerwise fashion, to approximate the local complexity in the prescribed neighborhood for a given layer.
Optimization Steps
@@
Refer to captionRefer to captionRefer to caption
Opt. Step 77035                 Opt. Step 83375                 Opt. Step 95381
Figure 7: Grokking visualized. We induce grokking by randomly initializing a $4$ depth $200$ width ReLU MLP and scaling the initialized parameters by eight following (Liu et al., 2022). In the leftmost figure, we can see that the grokking is visible for both the test samples as well as adversarial examples generated using the test set. We see that the network robustness, periodically increases. By visualization the partition and curvature of the function across a 2D slice of the input space (Humayun et al., 2023a), we see that the network periodically increases the concentration of non-linearity around its decision boundary, making the boundary sharper at each robustness peak. This occurs even when the network doesn’t undergo delayed generalization (Figure 2). As the local complexity around the decision boundary increases, the local complexity around data points farther from the decision boundary decreases (Figure 26).
@@
Figure 8: Region migration in modular addition. By measuring the local complexity for the GeLU activated fully connected layers of a Transformer architecture, we see that here as well, region migration occurs during grokking.
Accuracy
Local Complexity
Optimization Steps
Figure 9: Delayed robustness in LLMs. Grokking observed in a GPT architecture with 12 heads and 12 layers trained on next character prediction using the Shakespeare Text Dataset. We see that the second local complexity descent starts prior to the test acc. peak, and descent continues while the network groks ${\epsilon}=0.03$ $\ell_{\infty}$-PGD adversarial examples, generated in the token embedding space. Approximate input space partition visualized in Figure 19 and Figure 20.
Sensitivity of approximation to $P$ and $r$.
One of the possible limitations of local complexity measure is the deformation of the local neighborhood when its passed through a network from layer to layer, as shown in Figure 4. For different radius $r$ of the input space neighborhood ${\bm{V}}_{x}$ centered on any arbitrary data point $x$, we compute the graph eccentricity (Xu et al., 2021) of ${\bm{V}}_{x}$ after being embedded by different layers of a CNN. We present the results in Figure 5 for $1000$ different training data points for a CNN trained on CIFAR10. The higher the change of eccentricity compared to the input space (index 0), the more likely the neighborhood gets deformed, leading to less reliable approximation. We see that below a certain radius value, deformation by the CNN is limited and does not exponentially increase. In subsequent experiments however, e.g., Figure 27, we have observed that the dynamics of local complexity is similar between large and small $r$ neighborhoods. We present more validation experiments in Appendix A. Our proposed method can also be used to approximate the input space partition formed by a neural network. In Figure 25 we compare the partition approximated via LC computations on a grid, with analytically computed partition via (Humayun et al., 2023b). In Figure 19 and Figure 20 we present the input space partition approximated for a GPT architecture before and after delayed robustness occurs.
Experimental Setup.
@@ 3 Local Complexity Training Dynamics and Grokking
3.1 Emergence of a Robust Partition
@@
Figure 11: Token embedding space partition formed by the GeLU activated MLP of Block 12 of the GPT model mentioned in Figure 9. The partition is approximated by computing LC on a $512\times 512$ grid on a 2D subspace in the token embedding space. Note that after grokking (right), the three random samples used to determine the 2D subspace, has visibly lower local complexity in its immediate neighborhood.
@@
Local complexity as a progress measure.
@@
Relationship with Circuits.
@@
Table 1: Summary of all the experiments showing the relationship between delayed generalization and training/model hyperparameters.
Training Intervention Dataset Model Description Effect on Adv. Grokking
Increasing Batchsize CIFAR10 Resnet18 Batchsize Increased from 64 to 512 Expedited (Figure 17)
Increasing Parameters: Depth MNIST MLP Increasing Depth from 2 to 6 Expedited (Figure 27)
Increasing Parameters: Depth Shakespeare Text GPT Increasing number of layers from 6 to 12 Expedited
Increasing Parameters: Width MNIST MLP Increasing Width from 20 to 2000 Expedited (Figure 14)
Increasing Parameters: Width CIFAR10 Resnet18 Increasing Width from 16 filters to 64 filters Expedited
Increasing Regularization MNIST MLP Weight Decay increased from 0. to 1 Delayed (Figure 30)
Increasing Training Data MNIST MLP Training data increased from 1K to 60K Expedited (Figure 32)
Increasing Training Data (Randomly Labeled) MNIST MLP Training data increased from 1K to 60K Delayed (Figure 13)
No BatchNorm ->BatchNorm CIFAR10 Resnet18 Adding BatchNorm Layer after convolution Does not occur/significantly delayed
No BatchNorm ->BatchNorm CIFAR10 CNN Adding BatchNorm Layer after convolution Does not occur/significantly delayed
Refer to captionRefer to caption